# Document Reconstruction Unlocks Scalable Long-Context RLVR

### v2

@@ 5.1 Main Results
Figure 4: Average scores of RULER. We compare performance of dense and sparse rewards.
@@ 5.3 Robustness to Option Mixture Ratios
Figure 5: Average scores of RULER. We compare the performance of different option length mixture ratios.
@@ 5.4 Longer Documents Bring Consistent Improvement
Figure 6: We report the average score of RULER
@@ 5.7 Analysis of Option Length
In this part, we study the influence of sample difficulty on model performance. Intuitively, a sample with more candidate options is more challenging, as the model must search over a larger permutation space and rely more heavily on global contextual cues. Specifically, we randomly down-sample 5,000 documents from the training corpus and create four distinct training sets, where each set is constructed using a fixed option length $K\in\{2,4,6,8\}$, without mixing different values of $K$ within the same set. We train an independent model for each curated training set, where each model is associated with a specific $K$.
@@ 6 Related Work
Reinforcement Learning with Verifiable Rewards provides an objective and scalable framework for improving the reasoning capabilities of large language models by supervising them with ground-truth answers (Lambert et al., 2025; Guo et al., 2025). This paradigm has proven particularly effective in domains with unambiguous correctness criteria, including mathematics and programming. Previous work shows that RLVR can elevate models to expert-level reasoning performance by encouraging the discovery of correct internal reasoning trajectories through outcome-only rewards (OpenAI et al., 2024; Kim et al., 2025; Huang and Yang, 2025; Mo et al., 2025; Wang et al., 2026b). Most existing RLVR studies focus on self-contained reasoning tasks, where the central challenge is to recover a valid trajectory (Yue et al., 2025). In these settings, emergent behaviors, such as self-reflection, have been identified as important contributors to performance gains (Gandhi et al., 2025). Recent reasoning-oriented models, including OpenAI’s o1 series and DeepSeek, have popularized RLVR-based training pipelines and reinforcement learning algorithms such as GRPO.
@@
Training large language models for long-context reasoning poses challenges that differ fundamentally from those addressed by standard reinforcement learning with verifiable rewards. The outcome reward provides limited guidance in long-context settings, where success depends on identifying and grounding relevant evidence from extensive context (Wan et al., 2025). Most existing approaches to long-context training rely on supervised fine-tuning with synthetic data rather than reinforcement learning (Li et al., 2024; Yen et al., 2025; Chen et al., 2025b). Prior work pads questions with unrelated passages, shuffles document order, or fills contexts with irrelevant text to artificially increase sequence length (Trivedi et al., 2023). However, synthetic-data construction often introduces superficial complexity and fails to produce high-quality and challenging examples, limiting the effectiveness of the resulting long-context capabilities. Despite these limitations, improving long-context reasoning has become increasingly important due to the rapid emergence of agent applications (Zhao et al., 2024; Team et al., 2025a; b; Prabhakar et al., 2025; Gandhi et al., 2026). Thus, developing scalable and principled training methods that move beyond synthetic context expansion and provide effective supervision for long-context training is becoming urgent for advancing large language models.
@@ 7 Conclusion
In this work, we present an unsupervised reinforcement learning framework for improving the long-context capabilities of large language models through document reconstruction. By formulating reconstruction as a sequential decision-making problem with verifiable rewards derived directly from raw documents, our approach eliminates the need for manually curated long-context QA data or teacher-model supervision. This enables a scalable and principled alternative to existing long-context training paradigms.
Extensive experiments on RULER and LongBench v2 demonstrate that reconstruction-based RLVR effectively enhances long-context performance across multiple backbone models and context lengths. Beyond empirical gains, our findings highlight document structure itself as a valuable and underexplored supervision signal for long-context learning. We hope that this work motivates further research into unsupervised and self-supervised training objectives that leverage intrinsic structure in raw data and contributes to the development of more capable and scalable long-context language models.

### v3

@@ 5.1 Main Results
Model #Data RULER LongBench
Qwen2.5-7B-Instruct-1M
   SFT 46K 64.5 33.2
   Ours 14K 68.3 33.8
LLaMA-3.1-8B-Instruct
   SFT 46K 64.7 31.2
   Ours 14K 61.2 30.4
Table 1: Comparison between SFT and our reconstruction training on RULER (AVG) and LongBench v2 (Overall).
@@
Comparison to SFT
Chen et al. (2026) employ a powerful teacher model to curate 46K context-specific QA pairs for supervised fine-tuning. They report results on LongBench v2 and only the QA subset of RULER. In Table 1, we compare our results with theirs. With only 14K training samples and without relying on context-specific QA pairs, our approach achieves superior performance on Qwen2.5-7B-Instruct-1M, while slightly lags behind on LLaMA-3.1-8B-Instruct.
@@
Figure 4: Average scores on RULER. We compare the performance of dense and sparse rewards.
Figure 5: Average scores on RULER. We compare the performance of different option length mixture ratios.
@@ 5.4 Longer Documents Bring Consistent Improvement
Figure 6: We report the average score of RULER.
@@ 5.7 Analysis of Option Length
In this part, we study the influence of sample difficulty on model performance. We randomly sample 5,000 documents from the training corpus and create four distinct training sets, where each set is constructed using a fixed option length $K\in\{2,4,6,8\}$, without mixing different values of $K$ within the same set. We train an independent model for each curated training set, where each model is associated with a specific $K$.
@@ 6 Related Work
Reinforcement Learning with Verifiable Rewards provides an objective and scalable framework for improving the reasoning capabilities of large language models by supervising them with ground-truth answers (Lambert et al., 2025; Guo et al., 2025). Previous work shows that RLVR can elevate models to expert-level reasoning performance by encouraging the discovery of correct internal reasoning trajectories through outcome-only rewards (OpenAI et al., 2024; Kim et al., 2025; Huang and Yang, 2025; Mo et al., 2025; Wang et al., 2026b). Most existing RLVR studies focus on self-contained reasoning tasks, where the central challenge is to recover a valid trajectory (Yue et al., 2025). In these settings, emergent behaviors, such as self-reflection, have been identified as important contributors to performance gains (Gandhi et al., 2025).
@@
Training large language models for long-context reasoning poses challenges that differ fundamentally from those addressed by standard reinforcement learning with verifiable rewards. The outcome reward provides limited guidance in long-context settings, where success depends on identifying and grounding relevant evidence from extensive context (Wan et al., 2025). Most existing approaches to long-context training rely on supervised fine-tuning with synthetic data rather than reinforcement learning (Li et al., 2024; Yen et al., 2025; Chen et al., 2025b). Prior work pads questions with unrelated passages, shuffles document order, or fills contexts with irrelevant text to artificially increase sequence length (Trivedi et al., 2023). Improving long-context reasoning has become increasingly important due to the rapid emergence of agent applications (Zhao et al., 2024; Team et al., 2025a; b; Prabhakar et al., 2025; Gandhi et al., 2026).
@@ 7 Conclusion
In this work, we present an unsupervised reinforcement learning framework for improving the long-context capabilities of large language models. By formulating reconstruction as a sequential decision-making problem with verifiable rewards derived directly from raw documents, our approach eliminates the need for manually curated long-context data or teacher-model supervision. This enables a scalable and principled alternative to existing long-context training paradigms. Extensive experiments on RULER and LongBench v2 demonstrate that reconstruction-based RLVR effectively enhances long-context performance across multiple backbone models and context lengths. We hope that this work motivates further research into unsupervised and self-supervised training objectives that leverage intrinsic structure in raw data and contributes to the development of more capable and scalable long-context language models.